The College of Westchester ACC320 Managerial Mr. Paris Instructor Review for Final Exam Additional Problems 1. The Co. manufactures Zeon for its product. The Mfg. cost is $78 per batch including Fixed Costs of $22 per batch. A proposal is made from an outside source for $48 per Batch plus $10 freight. Prepare a differential analysis for the outside purchase proposal. 2. Product Zeon is normally sold for $97 per unit. A special price of $84 is offered for the export market. The variable production cost is $68 per unit. An additional export tari of 25% of the sales price must be paid on all export products. Calculate the differential ff income or loss per unit from selling Zeon for export. 3. A project has estimated annual cash flows of $71500 for 8 years and is estimated to cost $365400. Assume a minimum desired rate of return of 10%, using the present value of an Annuity of $1 at compound interest table, determine A) the net present value of the project and B) the present value index, rounded to two decimal places.
Answer #1
Make Buy Difference (Buy-Make)
Variable cost of manufacture 56 0 -56
Fixed cost of manufacture 22 22 0
Purchase 0 58 58
78 80 2
As the differential cost for Buy is $2 (higher by $2), the Co. should manufacture Zeon.
Variable cost of manufacture per unit 68
Net revenue from export, per unit = 84*(1-25%) = 63
Loss per unit -5
PV of annual cash inflows = 71500*5.3349 = $      3,81,445
Less: Initial investment $      3,65,400
NPV $          16,045
PI = 381445/365400 = 1.04